# Inference on multiple quantiles in regression models by a rank-score approach

### v1

@@ 5 Simulation study
For all simulation studies, we set the sample size to $n=100$ and the significance level to $\alpha=0.05$. For readability, the dependence on $n$ is suppressed in the mathematical notation. Covariate pairs $(x_{i},z_{i})^{\top}$ are drawn from a bivariate normal distribution with mean zero and an equicorrelated covariance matrix with correlation $0.3$, while different generating mechanisms are considered for the responses $y_{i}$. We define five individual hypotheses, each testing whether the effect of $x_{i}$ is zero at the quantile levels $\{0.1,0.25,0.5,0.75,0.9\}$. This results in $2^{5}-1=31$ possible intersection hypotheses, excluding the one corresponding to the empty set, as defined in Section 3. Each simulation scenario is replicated 1,0001{,}000 times.
First, we examine the type I error control of the Wald-type test of koenker2005quantile and the proposed rank-score test. In this setting, $y_{i}\sim\mathcal{N}(0.5\,z_{i},1)$, so there is no effect of $x_{i}$ at any quantile level, and all $31$ intersection hypotheses are tested. The resulting $p$-values, averaged over simulations, are shown in Fig. 2. Only three of the 31 $p$-values fall within the confidence bands for the Wald-type tests, indicating inadequate type I error control, whereas all rank-score $p$-values lie within the bands, showing proper control.
Figure 2: Type I error control for Wald-type and rank-score tests for intersection hypotheses. Average $p$-values; solid grey lines correspond to $\alpha$, and dotted lines represent the $95\%$ simulation confidence interval for the empirical type I error.
Subsequently, we assess and compare the power of the traditional Bonferroni correction and the proposed closed-testing procedure, both applied to rank-score tests. Three data-generating mechanisms are considered:
@@
(iii) $y_{i}\sim\mathcal{N}\left(0.5+\beta x_{i}+0.5\,z_{i},\,1+|x_{i}|\right).$
Here $t_{5}$ is the Student’s $t$-distribution with $5$ degrees of freedom, while $\mathrm{Exp}(\lambda_{i})$ is the exponential distribution with rate $\lambda_{i}$. The operator $\cdot]_{-30}^{30}$ denotes the value of its argument truncated to the interval [30,30][-30,30]. The parameter of interest, $\beta$, takes values in $\{0.2,0.4,0.6,0.8\}$ in the first two settings, and in $\{0.6,0.8,1.0,1.2\}$ in the third. For any quantile $\tau$, the true coefficient $\beta(\tau)$ depends not only on $\beta$, shared across quantiles, but also on quantile-specific features of the generating distribution. This follows from the definition of the quantile regression coefficient from Section 2.
Figure 3: Empirical power for individual hypotheses. Solid lines represent the closed testing procedure, and dotted lines represent the Bonferroni correction, both applied to rank-score tests.
The power behavior, shown in Fig. 3, is evaluated for each quantile and computed as the proportion of simulations in which an effect of $x_{i}$ is detected. The three scenarios reveal several common patterns. Overall, our procedure attains higher power than Bonferroni, with only minor exceptions, and ensures consistent convergence toward the maximum attainable power as the true parameter increases. When examining individual quantiles, power gains are most pronounced for the more challenging ones, corresponding to regions of the distribution with lower density. This pattern emerges in both tails for the first and third symmetric scenarios, and in the right tail for the second scenario, due to the skewness of the exponential distribution. Analogous results, not reported to avoid repetitions, were obtained for $\gamma=0$ and for correlations between $x_{i}$ and $z_{i}$ equal to 0 and $0.6$, indicating that the findings are robust to changes in nuisance parameters and covariate dependence.

### v2

@@ 6 Simulation study
For all simulation studies, we set the sample size to $n=100$ and the significance level to $\alpha=0.05$. For readability, the dependence on $n$ is suppressed in the mathematical notation. Covariate pairs $(x_{i},z_{i})^{T}$ are drawn from a bivariate normal distribution with mean zero and an equicorrelated covariance matrix with correlation $0.3$, while different generating mechanisms are considered for the responses $y_{i}$. Each simulation scenario is replicated 1,0001{,}000 times.
First, we define five individual hypotheses, each testing whether the effect of $x_{i}$ is zero at the quantile levels $\{0.1,0.25,0.5,0.75,0.9\}$. This results in $2^{5}-1=31$ possible intersection hypotheses, excluding the one corresponding to the empty set, as defined in Section 3. We then examine the type I error control of the Wald-type test of Koenker (2005) and the proposed generalized rank-score test, fixing $B=I_{5}$ (or the appropriate sub-matrix for each hypothesis) as in Section 5. In this setting, $y_{i}\sim\mathcal{N}(0.5+0.5\,z_{i},1)$, so there is no effect of $x_{i}$ at any quantile level. The resulting $p$-values, averaged over simulations, are shown in Fig. 2. Only three of the 31 $p$-values fall within the confidence bands for the Wald-type tests, indicating inadequate type I error control, whereas all generalized rank-score $p$-values lie within the bands, showing proper control. Power is then evaluated by computing the proportion of simulations in which an effect of $x_{i}$ is detected, and compared between the generalized rank-score tests with $B=A^{-1}$ (Section 4) and $B=I_{5}$ (Section 5), or appropriate sub-matrices based on each specific hypothesis. Under $y_{i}\sim\mathcal{N}(0.5+0.6x_{i}+0.5z_{i},1+|x_{i}|)$, Fig. 3 shows that the identity weighting improves power for most intersection hypotheses, with no gains for individual hypotheses along the bisector as expected.
Figure 2: Type I error control for Wald-type and generalized rank-score tests with $B=I_{5}$ for intersection hypotheses. Average $p$-values; solid grey lines correspond to $\alpha$, and dotted lines represent the $95\%$ simulation confidence interval for the empirical type I error.
Figure 3: Empirical power of the generalized rank-score tests for the $31$ intersection hypotheses, comparing the identity weighting ($B=I_{5}$) with the inverse-covariance weighting ($B=A^{-1}$).
Subsequently, we assess and compare the power of the Holm-Bonferroni and Bonferroni corrections with that of the proposed closed-testing procedure. Three data-generating mechanisms are considered:
@@
(iii) $y_{i}\sim\mathcal{N}\left(0.5+\beta x_{i}+{\gamma}z_{i},\,1+|x_{i}|\right).$
Here $t_{5}$ is the Student’s $t$-distribution with $5$ degrees of freedom, SN denotes the skew-normal distribution, and $\gamma$ is set to $0.5$.
For any quantile $\tau$, the true coefficient $\beta(\tau)$ depends not only on $\beta$, shared across quantiles, but also on quantile-specific features of the generating distribution. This follows from the definition of the quantile regression coefficient from Section 2.
Figure 4: Empirical power for individual hypotheses. Solid black lines represent the closed testing procedure, while dashed and dotted grey lines represent the Holm-Bonferroni and Bonferroni corrections, respectively, all applied to generalized rank-score tests with $B=f\circ F^{-1}(\tau)$, $B=I_{3}$ and $B=\text{diag}\{\Delta\}^{-1}$.
We compare the power of the proposed closed testing procedure with that of the Bonferroni and Holm–Bonferroni corrections for $\beta=0.6$, focusing on extreme quantile levels, namely $\{0.05,0.1,0.15\}$, and considering three different specifications of $B$. We emphasize that the choice of the weighting matrix must be made prior to observing the data in order to avoid cherry picking. Nevertheless, Fig. 4 shows that the identity specification represents a safe choice, even in challenging settings corresponding to regions of the distribution with low density. Analogous results, not reported to avoid repetitions, were obtained for $\gamma=0$ and $\beta\in\{0.4,0.8,1.2\}$, and for correlations between $x_{i}$ and $z_{i}$ equal to 0 and $0.6$, indicating that the findings are robust to changes in nuisance parameters and covariate dependence.
Figure 5: Empirical power for individual hypotheses. Solid black lines represent the closed testing procedure, while the dashed and dotted grey lines represent the Holm-Bonferroni and Bonferroni corrections, respectively, all applied to generalized rank-score tests with $B=I_{9}$.
Finally, the power behavior for $B=I_{9}$, shown in Fig. 5, is evaluated at each decile quantile. The three scenarios reveal several common patterns. Overall, our procedure attains higher power than traditional multiplicity corrections, with only minor exceptions, and displays consistent convergence toward the maximum attainable power as the true parameter increases. As before, analogous results (not reported) were obtained for different values of $\gamma$ and $\beta$, as well as for different correlations between $x_{i}$ and $z_{i}$.